OpenStudy (anonymous):

Is .63 the eccentricity of the ellipse x^2/49 + y^2/81 =1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to two decimal places

OpenStudy (anonymous):

Thank you. I'm having a hard time figuring this out. I know the formula is e = c^2/a but I don't get this answer. I'm guessing. Any suggestions?

OpenStudy (anonymous):

\[\sqrt{1-(\frac{9}{7})^2}\]

OpenStudy (anonymous):

or \[\sqrt{\frac{a^2-b^2}{a^2}}\]

OpenStudy (anonymous):

by which i mean \[\sqrt{\frac{a^2-b^2}{a^2}}=\sqrt{1-(\frac{b}{a})^2}\] so you can use whatever you find easiest

OpenStudy (anonymous):

sorry i had your backwards. it should have been \[\sqrt{1-(\frac{7}{9})^2}\]

OpenStudy (anonymous):

the thing inside the radical must be positive, so the bigger number goes irst or in the denominator

OpenStudy (anonymous):

My lesson says the formulat is y = a^2/c so I get 81/4 = 20.25. Where am I going wrong?